# A Framework for Objective-Driven Dynamical Stochastic Fields

### v1

@@ 2.2 Field Formulation
Imagine a system composed of multiple entities, each labeled by $x\in{\mathcal{X}}$. The locality principle implies that they each only interact with their neighbors. In the abstract formulation of our model, as introduced in Section 1, the ground space ${\mathcal{X}}$ is only required to have a topology that defines neighborhood relations. In this paper, we specifically model this space as a finite directed graph, leaving more general cases for future study.
@@
Therefore, the locality principle implies that the local dynamics of ${\omega}(t,x)$ only depends on its immediate neighborhood ${{\mathcal{U}}}_{x}$. Formally, it implies an independence relation generalized from equation 2.14.
@@
$\langle{\hat{\bm{e}}}_{{\omega}^{\prime}_{|{\Omega}_{x}}},{{\mathbf{G}}}(x){\hat{\bm{e}}}_{{\omega}_{|{{\mathcal{U}}}_{x}}}\rangle=\lim_{{\Delta t}\to 0}\frac{p_{{\Delta t}}({\omega}^{\prime}_{|{\Omega}_{x}}\mid{\omega}_{|{{\mathcal{U}}}_{x}})-\delta_{{\omega}_{|{\Omega}_{x}}}^{{\omega}^{\prime}_{|{\Omega}_{x}}}}{{\Delta t}}.$ (2.29)
@@
$\bar{{\mathbf{G}}}(x){\hat{\bm{e}}}_{{\omega}}=({{\mathbf{G}}}(x){\hat{\bm{e}}}_{{\omega}_{|{{\mathcal{U}}}_{x}}})\otimes{\hat{\bm{e}}}_{{\omega}_{|{\Omega}/{\Omega}_{x}}},$ (2.30)
@@
where the commutator $[{{\mathbf{G}}}(x),{{\mathbf{G}}}(x^{\prime})]={{\mathbf{G}}}(x){{\mathbf{G}}}(x^{\prime})-{{\mathbf{G}}}(x^{\prime}){{\mathbf{G}}}(x)$.
@@ 2.3 A Path Integral Formalism
We begin by considering the entire system and examining the probability that it starts in configuration ${\omega}_{0}={\omega}$ at time 00 and ends up in configuration ${\omega}_{T}={\omega}^{\prime}$ at time $T$. This propagation probability is given by:
$p_{t}({\omega}^{\prime}\mid{\omega})=\langle{\hat{\bm{e}}}_{{\omega}^{\prime}},e^{{{\mathbf{G}}}T}{\hat{\bm{e}}}_{{\omega}}\rangle.$ (2.33)
We can see that the system may go from ${\omega}$ to ${\omega}^{\prime}$ via multiple possible paths, each with different probability. Therefore, one may be curious to observe what happens in the middle and what path the system takes. To do this, let us try to observe the system every ${\Delta t}=T/N$ time for $N$ times. Suppose we observe a path
@@
$p({\omega}_{0:T,{\Delta t}}\mid{\omega}_{0})=\prod_{n=1}^{N}p_{\Delta t}({\omega}_{n{\Delta t}}\mid{\omega}_{(n-1){\Delta t}})=\prod_{n=1}^{N}\langle{\hat{\bm{e}}}_{{\omega}_{n{\Delta t}}},e^{{{\mathbf{G}}}{\Delta t}}{\hat{\bm{e}}}_{{\omega}_{(n-1){\Delta t}}}\rangle$ (2.35)
@@
$p_{T}({\omega}^{\prime}\mid{\omega})=\sum_{{\omega}_{0:T,{\Delta t}}}^{{\omega}\to{\omega}^{\prime}}p({\omega}_{0:T,{\Delta t}}\mid{\omega}_{0})=\sum_{{\omega}_{0:T,{\Delta t}}}^{{\omega}\to{\omega}^{\prime}}\prod_{n=1}^{N}\langle{\hat{\bm{e}}}_{{\omega}_{n{\Delta t}}},e^{{{\mathbf{G}}}{\Delta t}}{\hat{\bm{e}}}_{{\omega}_{(n-1){\Delta t}}}\rangle,$ (2.36)
@@
It turns out, with some algebraic manipulations, we can obtain something analogous to the Lagrangian of a physical system. However, unlike systems with smooth a configuration space, our system evolves in a discrete configuration set ${\Omega}$. To properly analyze its jump behavior, we require a specialized function—the unit impulse function, which is a formulation of the Dirac delta function. This should not be confused with the delta indicator $\delta_{\omega}^{{\omega}^{\prime}}$ that we have being using.
@@
We may also recognize this term as a modified log-likelihood, since $G^{{\omega}_{t}^{+}}_{{\omega}_{t}}$ represents the probability the system transitioning from ${\omega}_{t}$ to ${\omega}_{t}^{+}$.
@@
Notation. We begin by classifying different paths by the number of jumps, i.e., the number of configuration changes. Denote ${\omega}_{0:T,{\Delta t},k}$ as a path having $k$ jumps, and define $t_{n}=n{\Delta t}$. For any function of a path $f({\omega}_{0:T})$, adopting the conventional path integral notation, we denote the integration of $f$ over all paths going from ${\omega}$ to ${\omega}^{\prime}$ in time $T$ as the following.
@@
Having developed necessary machinery, a path integral formulation of the system can be derived as follows.
@@
$p_{T}({\omega}^{\prime}\mid{\omega})=\langle{\hat{\bm{e}}}_{{\omega}^{\prime}},e^{{{\mathbf{G}}}T}{\hat{\bm{e}}}_{{\omega}}\rangle=\int{\mathcal{D}}{\omega}\ \exp\left\{-\sum_{x}\int_{0}^{T}{\mathop{}\!\mathrm{d}t}\ L(w_{t,x},w_{t,x}^{+})\right\},$ (2.43)
@@ 2.4 Completing the Generator Formalism and the Deterministic Limit
To introduce the temporal characterization to the generator formalism, let us begin by motivating it through the following interesting problem. Suppose the system has a fixed “energy” such that the frequency of its jump between configurations is constant. Formally, noting that $|G^{{\omega}}_{{\omega}}|$ is the rate at which the system jumps from ${\omega}$ to other configurations, let us demand
@@
Observe that an impulse essentially contributes a value each time a jump occurs. However, this behavior cannot be captured directly by our space ${\mathcal{H}}$. Therefore, we need to expand it to include the temporal neighborhood into consideration, resulting in a larger space ~${\widetilde{\mathcal{H}}}({\Omega})$.
Definition 2.19 ( Space ~${\widetilde{\mathcal{H}}}({\Omega})$).
@@
~${\widetilde{\mathcal{H}}}({\Omega}):={\mathcal{H}}({\Omega}\times{\Omega})\cong{\mathcal{H}}({\Omega})\otimes{\mathcal{H}}({\Omega}).$ (2.47)
A basis set of ~${\widetilde{\mathcal{H}}}({\Omega})$ is denoted as
${\hat{\bm{e}}}_{{\omega}^{\prime}{\omega}}\quad\text{for all }\ {\omega}^{\prime},{\omega}\in{\Omega}.$ (2.48)
When the context is clear, we only use ${\widetilde{\mathcal{H}}}$ to denote ~${\widetilde{\mathcal{H}}}({\Omega})$.
Then, in order to investigate equation 2.45, we need to lift the generator ${{\mathbf{G}}}\colon{\mathcal{H}}\to{\mathcal{H}}$ to ${\widetilde{\mathbf{G}}}\colon{\mathcal{H}}\to{\widetilde{\mathcal{H}}}$, as follows.
Notation. We denote ${{\widetilde{\mathbf{A}}}}_{{\omega}^{\prime}}^{{\omega}}\colon{\mathcal{H}}\to{\widetilde{\mathcal{H}}}$ as a linear operator that transforms ${\hat{\bm{e}}}_{{\omega}}$ to ${\hat{\bm{e}}}_{{\omega}^{\prime}{\omega}}$. Formally
${{\widetilde{\mathbf{A}}}}_{{\omega}^{\prime}}^{{\omega}}{\hat{\bm{e}}}_{{\omega}^{\prime\prime}}=\begin{cases}{\hat{\bm{e}}}_{{\omega}^{\prime}{\omega}},\qquad&\text{if}\ {\omega}={\omega}^{\prime\prime}\\ 0,&\text{otherwise}.\end{cases}$ (2.49)
Definition 2.20 ( Lift of the Generator ${\widetilde{\mathbf{G}}}\colon{\mathcal{H}}\to{\widetilde{\mathcal{H}}}$).
The lift of generator ${\widetilde{\mathbf{G}}}\colon{\mathcal{H}}\to{\widetilde{\mathcal{H}}}$ is defined to be
${\widetilde{\mathbf{G}}}{\hat{\bm{e}}}_{\omega}=G^{{\omega}^{\prime}}_{{\omega}}{\hat{\bm{e}}}_{{\omega}^{\prime}{\omega}},\qquad\text{or equivalently}\qquad{\widetilde{\mathbf{G}}}=G^{{\omega}^{\prime}}_{{\omega}}{{\widetilde{\mathbf{A}}}}_{{\omega}^{\prime}}^{{\omega}}.$ (2.50)
@@
It corresponds to an operator ${\bm{\Gamma}}:{\widetilde{\mathcal{H}}}\to{\mathbb{R}}$ defined as
${\bm{\Gamma}}:=\gamma_{{\omega}^{\prime}{\omega}}{{\check{\bm{\eta}}}}^{{\omega}^{\prime}{\omega}},\qquad\text{where }\ \forall{\omega}:\gamma_{{\omega}{\omega}}=0.$ (2.52)
@@
$\mathbb{E}\left[\int_{0}^{T}{\mathop{}\!\mathrm{d}t}\ \gamma(t,{\omega},{\omega}^{+})\Bigm{|}{\omega}_{0}={\omega}\right]=\int_{0}^{T}{\mathop{}\!\mathrm{d}t}\ {\bm{\Gamma}}{\widetilde{\mathbf{G}}}e^{{{\mathbf{G}}}t}{\hat{\bm{e}}}_{{\omega}}.$ (2.53)
@@
$\tilde{\bm{L}}\ {\hat{\bm{e}}}_{{\omega},{\omega}^{\prime}}=\begin{cases}-\log G^{{\omega}^{\prime}}_{{\omega}},\qquad&\text{if }\ {\omega}^{\prime}\neq{\omega},\\ 0,&\text{otherwise}.\end{cases}$ (2.54)
Note that if ${\omega}^{\prime}\neq{\omega}$ then $G^{{\omega}^{\prime}}_{{\omega}}\geq 0$ by definition. If $G^{{\omega}^{\prime}}_{{\omega}}=0$, we allow $\log 0=-\infty$ with the convention that $00\log 0=$ 0. By the above proposition, we can see that
$\mathbb{E}\left[\int_{0}^{T}{\mathop{}\!\mathrm{d}t}\ L(w_{t},w_{t}^{+})\Bigm{|}{\omega}_{0}={\omega}\right]=KT+\int_{0}^{T}{\mathop{}\!\mathrm{d}t}\ \tilde{\bm{L}}{\widetilde{\mathbf{G}}}e^{Gt}{\hat{\bm{e}}}_{{\omega}}.$ (2.55)
@@
$\tilde{\bm{L}}{\widetilde{\mathbf{G}}}{\hat{\bm{e}}}_{{\omega}}=-K\log K+H_{\omega}.$ (2.57)
@@ 3 Objective-Driven Dynamical Stochastic Field
From a perspective of artificial intelligence, such as solving machine learning tasks, it involves two aspects. The first is how the field evolves to minimize its objective value. The second is how the objective value is designed such that its minimization leads to desired behaviors. We discuss both aspects in this section, after introducing the necessary groundwork. We follow a similar approach to the previous section, beginning with a study of the entire system, then moving to the two-entity view, and finally arriving at the field formulation.
@@
The objective operator is a linear operator ${\bm{\Gamma}}\colon{\widetilde{\mathcal{H}}}\to{\mathbb{R}}$ defined by
${\bm{\Gamma}}:=\gamma_{{\omega}^{\prime}{\omega}}{{\check{\bm{\eta}}}}^{{\omega}^{\prime}{\omega}},\qquad\text{where }\ \forall{\omega}:\gamma_{{\omega}{\omega}}=0.$ (3.1)
@@
$\int_{0}^{T}{\mathop{}\!\mathrm{d}t}\ {\bm{\Gamma}}{\widetilde{\mathbf{G}}}{\hat{\bm{\varphi}}}(t)=\int_{0}^{T}{\mathop{}\!\mathrm{d}t}\ {\bm{\Gamma}}{\widetilde{\mathbf{G}}}e^{{{\mathbf{G}}}t}{\hat{\bm{\varphi}}}(0)$ (3.2)
is the expected value the system receives within time $T$ starting from state ${\hat{\bm{\varphi}}}(0)$. However, we want this objective value to characterize a fundamental property of the system, and it should depend on a minimal set of variables. Therefore, we define the objective value as a time average over an infinite time horizon.
Definition 3.2 (The Averaged State ${\hat{\bar{\bm{\varphi}}}}$ and the Objective Value ${\bar{\gamma}}$).
@@
${\hat{\bar{\bm{\varphi}}}}:=\lim_{T\to\infty}\frac{1}{T}\int_{0}^{T}{\mathop{}\!\mathrm{d}t}\ {\hat{\bm{\varphi}}}(t).$ (3.3)
@@
${\bar{\gamma}}:=\lim_{T\to\infty}\frac{1}{T}\int_{0}^{T}{\mathop{}\!\mathrm{d}t}\ {\bm{\Gamma}}{\widetilde{\mathbf{G}}}{\hat{\bm{\varphi}}}(t).$ (3.4)
The starting state ${\hat{\bm{\varphi}}}(0)$ will be either clear from context or irrelevant.
One might wonder if the above definitions are valid in the first place. In fact, their validness comes straightforwardly from the following result.
@@
${\hat{\bm{\varphi}}}(\infty):=\lim_{t\to\infty}{\hat{\bm{\varphi}}}(t)$ (3.5)
@@
Corollary 3.4 (${\hat{\bar{\bm{\varphi}}}}$ and ${\bar{\gamma}}$ Always Exist).
Given that the configuration set ${\Omega}$ is finite, the averaged state ${\hat{\bar{\bm{\varphi}}}}$ and the objective value ${\bar{\gamma}}$ always exist. Moreover, we have
${\hat{\bar{\bm{\varphi}}}}={\hat{\bm{\varphi}}}(\infty)\qquad\text{and}\qquad{\bar{\gamma}}={\bm{\Gamma}}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}.$ (3.6)
Having the necessary definitions, we formulate the principle of purposefulness as follow.
@@
Additionally, in the rest of the section, we will need the notion of ergodicity for the ease of analyses.
@@
The system is ergodic if ${\hat{\bar{\bm{\varphi}}}}$ does not depend on the initial state ${\hat{\bm{\varphi}}}(0)$.
The above property alleviates the need for excessive technical considerations in the remainder of this section. It ensures the well-definedness of many operations we will perform while maintaining a clear presentation. However, it is conjectured that only a weaker version of this property is necessary. For example, a straightforward relaxation would only require ${\bar{\gamma}}={\bm{\Gamma}}{\widetilde{\mathbf{G}}}{\hat{\bm{\varphi}}}(\infty)$ to be independent of the initial state. This is left for future work, and we are now prepared for a deeper investigation into the minimization of the objective value.

### v2

@@ 2.2 Field Formulation
Imagine a system composed of multiple entities, each labeled by $x\in{\mathcal{X}}$. The locality principle implies that each entity interacts only with its neighbors. In the abstract formulation of our model, as introduced in Section 1, the ground space $\mathcal{X}$} is only required to have a topology that defines neighborhood relations. In this paper, we specifically model this space as a finite directed graph, leaving more general cases for future study.
@@
Therefore, the locality principle implies that the local dynamics of ${\omega}(t,x)$ only depend on its immediate neighborhood ${{\mathcal{U}}}_{x}$. Formally, it implies an independence relation generalized from equation 2.12.
@@
$\langle{\omega}^{\prime}_{|{\Omega}_{x}}\rvert{{\mathbf{G}}}(x)\lvert{\omega}_{|{{\mathcal{U}}}_{x}}\rangle=\lim_{{\Delta t}\to 0}\frac{p_{{\Delta t}}({\omega}^{\prime}_{|{\Omega}_{x}}\mid{\omega}_{|{{\mathcal{U}}}_{x}})-\delta_{{\omega}_{|{\Omega}_{x}}}^{{\omega}^{\prime}_{|{\Omega}_{x}}}}{{\Delta t}}.$ (2.27)
@@
$\bar{{\mathbf{G}}}(x)\lvert{\omega}\rangle=({{\mathbf{G}}}(x)\lvert{\omega}_{|{{\mathcal{U}}}_{x}}\rangle)\otimes\lvert{\omega}_{|{\Omega}/{\Omega}_{x}}\rangle,$ (2.28)
@@
where the commutator ${{\mathbf{G}}}(x),{{\mathbf{G}}}(x^{\prime})]={{\mathbf{G}}}(x){{\mathbf{G}}}(x^{\prime})-{{\mathbf{G}}}(x^{\prime}){{\mathbf{G}}}(x)$.
@@ 2.3 A Path Integral Formalism
We begin by considering the entire system and examining the probability that it starts in configuration ${\omega}_{0}={\omega}$ at time 0 and ends up in configuration ${\omega}_{T}={\omega}^{\prime}$ at time $T$. This propagation probability is given by:
$p_{T}({\omega}^{\prime}\mid{\omega})=\langle{\omega}^{\prime}\rvert e^{{{\mathbf{G}}}T}\lvert{\omega}\rangle.$ (2.31)
We can see that the system may go from $\omega$} to $\omega$$^{\prime}$ via multiple possible paths, each with a different probability. Therefore, one may be curious to observe what happens in the middle and what path the system takes. To do this, let us try to observe the system every ${\Delta t}=T/N$ time for $N$ times. Suppose we observe a path
@@
$p({\omega}_{0:T,{\Delta t}}\mid{\omega}_{0})=\prod_{n=1}^{N}p_{\Delta t}({\omega}_{n{\Delta t}}\mid{\omega}_{(n-1){\Delta t}})=\prod_{n=1}^{N}\langle{\omega}_{n{\Delta t}}\rvert e^{{{\mathbf{G}}}{\Delta t}}\lvert{\omega}_{(n-1){\Delta t}}\rangle$ (2.33)
@@
$p_{T}({\omega}^{\prime}\mid{\omega})=\sum_{{\omega}_{0:T,{\Delta t}}}^{{\omega}\to{\omega}^{\prime}}p({\omega}_{0:T,{\Delta t}}\mid{\omega}_{0})=\sum_{{\omega}_{0:T,{\Delta t}}}^{{\omega}\to{\omega}^{\prime}}\prod_{n=1}^{N}\langle{\omega}_{n{\Delta t}}\rvert e^{{{\mathbf{G}}}{\Delta t}}\lvert{\omega}_{(n-1){\Delta t}}\rangle,$ (2.34)
@@
It turns out that, with some algebraic manipulations, we can obtain something analogous to the Lagrangian of a physical system. However, unlike systems with a smooth configuration space, our system evolves in a discrete configuration set $\Omega$}. To properly analyze its jump behavior, we require a specialized function, the unit impulse function, which is a formulation of the Dirac delta function. This should not be confused with the delta indicator $\delta_{\omega}^{{\omega}^{\prime}}$ that we have been using.
@@
We may also recognize this term as a modified log-likelihood, since $G^{{\omega}_{t}^{+}}_{{\omega}_{t}}$ represents the probability of the system transitioning from ${\omega}_{t}$ to ${\omega}_{t}^{+}$.
@@
Notation. We begin by classifying different paths by the number of jumps, i.e., the number of configuration changes. Denote ${\omega}_{0:T,{\Delta t},k}$ as a path having $k$ jumps, and define $t_{n}=n{\Delta t}$. For any function of a path $f({\omega}_{0:T})$, adopting the conventional path integral notation, we denote the integration of $f$ over all paths going from $\omega$} to $\omega$$^{\prime}$ in time $T$ as follows.
@@
Having developed the necessary machinery, a path integral formulation of the system can be derived as follows.
@@
$p_{T}({\omega}^{\prime}\mid{\omega})=\langle{\omega}^{\prime}\rvert e^{{{\mathbf{G}}}T}\lvert{\omega}\rangle=\int{\mathcal{D}}{\omega}\ \exp\left\{-\sum_{x}\int_{0}^{T}{\mathop{}\!\mathrm{d}t}\ L(w_{t,x},w_{t,x}^{+})\right\},$ (2.41)
@@ 2.4 Completing the Generator Formalism and the Deterministic Limit
To introduce the temporal characterization to the generator formalism, let us begin by motivating it through the following interesting problem. Suppose the system has a fixed “energy” such that the frequency of its jumps between configurations is constant. Formally, noting that $|G^{{\omega}}_{{\omega}}|$ is the rate at which the system jumps from $\omega$} to other configurations, let us demand
@@
Observe that an impulse essentially contributes a value each time a jump occurs. However, this behavior cannot be captured directly by our space $\mathcal{H}$}. Therefore, we need to expand it to include the temporal neighborhood into consideration, resulting in a larger space ${\widetilde{\mathcal{H}}}({\Omega})$.
Definition 2.19 ( Space ${\widetilde{\mathcal{H}}}({\Omega})$).
@@
${\widetilde{\mathcal{H}}}({\Omega}):={\mathcal{H}}({\Omega}\times{\Omega})\cong{\mathcal{H}}({\Omega})\otimes{\mathcal{H}}({\Omega}).$ (2.45)
A basis set of ${\widetilde{\mathcal{H}}}({\Omega})$ is denoted as
$\lvert{\omega}^{\prime}{\omega}\rangle\quad\text{for all }\ {\omega}^{\prime},{\omega}\in{\Omega}.$ (2.46)
When the context is clear, we only use $\widetilde{\mathcal{H}}$} to denote ${\widetilde{\mathcal{H}}}({\Omega})$.
[…]
Notation. We denote ${{\widetilde{\mathbf{A}}}}_{{\omega}^{\prime}}^{{\omega}}\colon{\mathcal{H}}\to{\widetilde{\mathcal{H}}}$ as a linear operator that transforms $\lvert{\omega}\rangle$ to $\lvert{\omega}^{\prime}{\omega}\rangle$. Formally
${{\widetilde{\mathbf{A}}}}_{{\omega}^{\prime}}^{{\omega}}\lvert{\omega}^{\prime\prime}\rangle=\begin{cases}\lvert{\omega}^{\prime}{\omega}\rangle,\qquad&\text{if}\ {\omega}={\omega}^{\prime\prime}\\ 0,&\text{otherwise}.\end{cases}$ (2.47)
[…]
${\widetilde{\mathbf{G}}}\lvert{\omega}\rangle=\sum_{{\omega}^{\prime}\in{\Omega}}G^{{\omega}^{\prime}}_{{\omega}}\lvert{\omega}^{\prime}{\omega}\rangle,\qquad\text{or equivalently}\qquad{\widetilde{\mathbf{G}}}=\sum_{{\omega},{\omega}^{\prime}\in{\Omega}}G^{{\omega}^{\prime}}_{{\omega}}{{\widetilde{\mathbf{A}}}}_{{\omega}^{\prime}}^{{\omega}}.$ (2.48)
@@
It corresponds to an operator $\bm{\Gamma}$$\widetilde{\mathcal{H}}$$\to{\mathbb{R}}$ defined as
${\bm{\Gamma}}:=\sum_{{\omega},{\omega}^{\prime}\in{\Omega}}\gamma_{{\omega}^{\prime}{\omega}}\langle{\omega}^{\prime}{\omega}\rvert,\qquad\text{where }\ \forall{\omega}:\gamma_{{\omega}{\omega}}=0.$ (2.50)
@@
$\mathbb{E}\left[\int_{0}^{T}{\mathop{}\!\mathrm{d}t}\ \gamma(t,{\omega},{\omega}^{+})\Bigm|{\omega}_{0}={\omega}\right]=\int_{0}^{T}{\mathop{}\!\mathrm{d}t}\ {\bm{\Gamma}}{\widetilde{\mathbf{G}}}e^{{{\mathbf{G}}}t}\lvert{\omega}\rangle.$ (2.51)
@@
$\tilde{\bm{L}}\ \lvert{\omega},{\omega}^{\prime}\rangle=\begin{cases}-\log G^{{\omega}^{\prime}}_{{\omega}},\qquad&\text{if }\ {\omega}^{\prime}\neq{\omega},\\ 0,&\text{otherwise}.\end{cases}$ (2.52)
Note that if ${\omega}^{\prime}\neq{\omega}$ then $G^{{\omega}^{\prime}}_{{\omega}}\geq 0$ by definition. If $G^{{\omega}^{\prime}}_{{\omega}}=0$, we allow $\log 0=$-$\infty$ with the convention that $0\log 0=0$. By the above proposition, we can see that
$\mathbb{E}\left[\int_{0}^{T}{\mathop{}\!\mathrm{d}t}\ L(w_{t},w_{t}^{+})\Bigm|{\omega}_{0}={\omega}\right]=KT+\int_{0}^{T}{\mathop{}\!\mathrm{d}t}\ \tilde{\bm{L}}{\widetilde{\mathbf{G}}}e^{Gt}\lvert{\omega}\rangle.$ (2.53)
@@
$\tilde{\bm{L}}{\widetilde{\mathbf{G}}}\lvert{\omega}\rangle=-K\log K+H_{\omega}.$ (2.55)
@@ 3 Objective-Driven Dynamical Stochastic Field
From a perspective of artificial intelligence, such as solving machine learning tasks, it involves two aspects. The first is how the field evolves to minimize its objective value. The second is how the objective value is designed such that its minimization leads to the desired behaviors. We discuss both aspects in this section, after introducing the necessary groundwork. We follow a similar approach to the previous section, beginning with a study of the entire system, then moving to the two-entity view, and finally arriving at the field formulation.
@@
[…]
${\bm{\Gamma}}:=\sum_{{\omega},{\omega}^{\prime}\in{\Omega}}\gamma_{{\omega}^{\prime}{\omega}}\langle{\omega}^{\prime}{\omega}\rvert,\qquad\text{where }\ \forall{\omega}:\gamma_{{\omega}{\omega}}=0.$ (3.1)
@@
$\int_{0}^{T}{\mathop{}\!\mathrm{d}t}\ {\bm{\Gamma}}{\widetilde{\mathbf{G}}}\lvert\varphi(t)\rangle=\int_{0}^{T}{\mathop{}\!\mathrm{d}t}\ {\bm{\Gamma}}{\widetilde{\mathbf{G}}}e^{{{\mathbf{G}}}t}\lvert\varphi(0)\rangle$ (3.2)
is the expected value the system receives within time $T$ starting from state $\lvert\varphi(0)\rangle$. However, we want this objective value to characterize a fundamental property of the system, and it should depend on a minimal set of variables. Therefore, we define the objective value as a time average over an infinite time horizon.
Definition 3.2 (The Averaged State $\lvert\bar{\varphi}\rangle$} and the Objective Value $\bar{\gamma}$}).
@@
${\lvert\bar{\varphi}\rangle}:=\lim_{T\to\infty}\frac{1}{T}\int_{0}^{T}{\mathop{}\!\mathrm{d}t}\ \lvert\varphi(t)\rangle.$ (3.3)
@@
${\bar{\gamma}}:=\lim_{T\to\infty}\frac{1}{T}\int_{0}^{T}{\mathop{}\!\mathrm{d}t}\ {\bm{\Gamma}}{\widetilde{\mathbf{G}}}\lvert\varphi(t)\rangle.$ (3.4)
The starting state $\lvert\varphi(0)\rangle$ will be either clear from context or irrelevant.
One might wonder if the above definitions are valid in the first place. In fact, their validity comes straightforwardly from the following result.
@@
$\lvert\varphi(\infty)\rangle:=\lim_{t\to\infty}\lvert\varphi(t)\rangle$ (3.5)
@@
Corollary 3.4 ($\lvert\bar{\varphi}\rangle$} and $\bar{\gamma}$} Always Exist).
Given that the configuration set $\Omega$} is finite, the averaged state $\lvert\bar{\varphi}\rangle$} and the objective value $\bar{\gamma}$} always exist. Moreover, we have
${\lvert\bar{\varphi}\rangle}=\lvert\varphi(\infty)\rangle\qquad\text{and}\qquad{\bar{\gamma}}={\bm{\Gamma}}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}.$ (3.6)
Having the necessary definitions, we formulate the principle of purposefulness as follows.
@@
Additionally, in the rest of this section, we will need the notion of ergodicity for the ease of analysis.
@@
The system is ergodic if $\lvert\bar{\varphi}\rangle$} does not depend on the initial state $\lvert\varphi(0)\rangle$.
The above property alleviates the need for excessive technical considerations in the remainder of this section. It ensures the well-definedness of many operations we will perform while maintaining a clear presentation. However, it is conjectured that only a weaker version of this property is necessary. For example, a straightforward relaxation would only require ${\bar{\gamma}}={\bm{\Gamma}}{\widetilde{\mathbf{G}}}\lvert\varphi(\infty)\rangle$ to be independent of the initial state. This is left for future work, and we are now prepared for a deeper investigation into the minimization of the objective value.